OpenStudy (anonymous):

x + 2y = -9 4x - 2y = 14 A. (-5,1) B. (-4,1) C. (1,-5) D. (1,-4)

OpenStudy (anonymous):

@Brill @Decarr432 @foolofatook @mthompson440

OpenStudy (calculusxy):

this seems like a system of equations problem... is there a certain way (substitution or elimination) that you are asked to solve this problem?

OpenStudy (anonymous):

idk @calculusxy

OpenStudy (calculusxy):

okay. so i think that using elimination will be much better. do you know how to solve systems using the elimination method?

OpenStudy (anonymous):

not realy

OpenStudy (calculusxy):

the elimination method means that we get rid of one variable (x or y) and then solve for the other variable that is present through zero pairs. for instance, if we have the equations \(x+3y=-14\) and \(-4x-3y=2\), we can get rid of the y variable because \(3y - 3y = 0y \text{ } \text{or simply 0}\).

OpenStudy (anonymous):

o got -4,1

OpenStudy (calculusxy):

then we would have the x variable to solve for \(x + (-4x) = -4x\) and \(-14 + 2 = -12\) \(-3x=-12\) \(x=4\)

OpenStudy (calculusxy):

sorry \(x + (-4x) = -3x\)

OpenStudy (anonymous):

ohh

OpenStudy (calculusxy):

similarly, in this situation \(x + 2y = -9\) \(4x - 2y = 14\) we can easily get rid of y because \(2y + (-2y) = 0y \text{ } \text{or 0}\) to solve for x we can have \(x + 4x = 5x\) and \(-9+14=5\) \(5x=5\) \(x=\text{?}\)

OpenStudy (calculusxy):

do you understand ?

OpenStudy (anonymous):

wount x be be 0

OpenStudy (calculusxy):

can you show me your work?

OpenStudy (anonymous):

well i dont really know how that stuffs works u know so used to saying it out load

OpenStudy (calculusxy):

do you understand the work that i showed?

OpenStudy (anonymous):

yeah man thank u

OpenStudy (calculusxy):

okay and remember to plug back in that x = 1 to x + 2y = -9 or the other equation and solve for y!